A body falls From a height of 200m
if gravitional ceases after 2 sec find the further time required to reach grounf

Grade:9

2 Answers

vikas askiitian expert
13 years ago

after 2 seconds velocity is v

v = u+at

u=0 ,a=-g

v = -2g =-20m/s               (downward)

distance covered in 2 second = S

 v2 = u2+2aS

u=0 , a=-g

S = v2/2a = 400/20 =20

it has covered 20 m so remaining distance from ground is 200-20 = 180m

now gravity there is no gravity effect so a=0

 now we use speed = distance/time

      t = d/v

   distance = 180 ,velocity = 20

    t = 180/20 = 9sec

thus body will take further 9 sec to reach the ground...

Fawz Naim
13 years ago

After travelling for 2sec the gravity ceases then the body will never reach the ground as there is no force pulling it down